The line, among the following, that touches the parabola y2
 = 4ax is [EAMCET 2005] 
 1) 3 x my am 0 ++ = 2) 2 x my am 0 −+ = 3) 2 x my am 0 + − = 4) 2 y mx am 0 ++ =

Grade:12th pass

1 Answers

ROSHAN MUJEEB
3 years ago
DEAR STUDENT
We can write lx + my + n = 0 in the form of y = Mx + c

So, my = -lx - n
y = -lx/m - n/m

Comparing with y = Mx + c,
M = -l/m
c = -n/m

We know that c = a/M
Mc = a

Putting the values of M and c,
-l/m x - n/m = a
ln/m^2 = a
ln = am^2

Hope This Helps You!